### v1

# RxRx3-core: Benchmarking drug-target interactions in High-Content Microscopy

Abstract
High Content Screening (HCS) microscopy datasets have transformed the ability to profile cellular responses to genetic and chemical perturbations, enabling cell-based inference of drug-target interactions (DTI). However, the adoption of representation learning methods for HCS data has been hindered by the lack of accessible datasets and robust benchmarks. To address this gap, we present RxRx3-core, a curated and compressed subset of the RxRx3 dataset, and an associated DTI benchmarking task. At just 18GB, RxRx3-core significantly reduces the size barrier associated with large-scale HCS datasets while preserving critical data necessary for benchmarking representation learning models against a zero-shot DTI prediction task. RxRx3-core includes 222,601 microscopy images spanning 736 CRISPR knockouts and 1,674 compounds at 8 concentrations. RxRx3-core is available on HuggingFace and Polaris, along with pre-trained embeddings and benchmarking code, ensuring accessibility for the research community. By providing a compact dataset and robust benchmarks, we aim to accelerate innovation in representation learning methods for HCS data and support the discovery of novel biological insights.
Dataset: https://huggingface.co/datasets/recursionpharma/rxrx3-core
Polaris Platform: https://polarishub.io/datasets/recursion/rxrx3-core
Benchmarks: https://github.com/recursionpharma/EFAAR_benchmarking
For more information about RxRx3-core please visit rxrx.ai/rxrx3-core
1 Introduction
Understanding and quantifying how cells respond to genetic and chemical perturbations — and relating these responses — is a central challenge in biological research  (Przybyla & Gilbert, 2022; Vincent et al., 2022). Advances in automated imaging platforms have transformed our ability to investigate cellular phenotypes triggered by a wide range of perturbations  (Boutros et al., 2015). High-Content Screening (HCS) systems, which integrate automated microscopy with robotic liquid handling, enable systematic profiling of cellular responses on an unprecedented scale. Publicly available HCS datasets, such as RxRx3  (Fay et al., 2023) and JUMP  (Chandrasekaran et al., 2023), exemplify this scalability, including millions of cellular images spanning hundreds of thousands of unique perturbations.
Figure 1: Overview of RxRx3-core, a subset of RxRx3 (Fay et al., 2023), designed for benchmarking representation learning models.
While the vast scope of HCS datasets creates opportunities to uncover novel biological insights, it also introduces technical challenges in deriving meaningful and interpretable features from these images. Traditionally, HCS images are processed using pipelines that rely on cell segmentation, feature extraction, and subsequent analyses tailored to specific datasets  (Caicedo et al., 2017). Although these workflows have enabled significant discoveries  (Boutros et al., 2015), they often require complex and resource-intensive optimization, through proprietary software or open-source tools  (Carpenter et al., 2006; Stirling et al., 2021). More recently, representation learning methods leveraging deep learning to learn feature representations directly from pixel values in HCS images have emerged as a promising alternative.
Despite their potential, the progress of representation learning in HCS has been limited by the scarcity of accessible and standardized datasets and benchmarks. Existing benchmarks largely focus on straightforward classification tasks, such as identifying cell cycle stages (Eulenberg et al., 2017), protein localization patterns (Ouyang et al., 2019), or mechanisms of action (Ljosa et al., 2013). However, these tasks fail to represent the complexity of genome-wide relationships that researchers often need to explore. Although recent large-scale datasets like RxRx3 and JUMP became available in 2023, their sheer size — over 100 TB each — along with the lack of standardized benchmarks, has impeded their widespread adoption.
Previous efforts to release smaller datasets specifically for benchmarking zero-shot prediction of drug-target interactions (DTI) based on compound target annotations, like CPJUMP1 (Chandrasekaran et al., 2024), have suffered from plate layout confounders, since well positions were not randomized between technical replicates. More recently, supervised learning has been applied to the CPJUMP1 dataset by treating the DTI task as a binary classification of gene-compound pairs (Iyer et al., 2024). However this approach was limited by the small number of perturbations in the dataset (302 compounds and 160 genes) and struggled to generalize to unseen genes. The Motive dataset (Arevalo et al., 2024) paired CellProfiler features extracted from the JUMP Cell Painting dataset (Chandrasekaran et al., 2023) covering 11,000 genes and 3,600 compounds with annotations aggregated from several public databases. This task, however, is framed as a graph learning task leveraging the extracted features rather than a benchmark designed to assess microscopy representation learning models. Benchmarking DTI prediction as a link prediction task introduces several issues including introducing false positives for annotated interactions in which one or both perturbations do not induce a cellular phenotype and the requirement to carefully design train/test splits to ensure model generalization is accurately assessed. Evaluating the proposed benchmarking task with an alternative model to CellProfiler is burdensome as it would require downloading and running inference on $>$15TB of imaging data and subsequently evaluating graph learning models on these features.
To address these limitations, we introduce RxRx3-core, a dataset specifically designed to facilitate research in representation learning for HCS data. At only 18 GB (17.5GB of images and 0.5Gb of pre-computed embeddings), RxRx3-core is significantly more accessible than its predecessors and is accompanied by benchmarks evaluating perturbation signal magnitude and zero-shot DTI prediction. By providing both a manageable dataset and well-defined benchmarks, we aim to accelerate innovation in representation learning methods for high-content phenotypic screening.
2 RxRx3-core Dataset
RxRx3-core is a compressed and accessible subset of RxRx3 (Fay et al., 2023), a dataset initially released by Recursion in 2023. RxRx3 contains 6-channel fluorescent microscopy images of HUVEC cells stained using a modified Cell Painting protocol  (Bray et al., 2016). The full dataset comprises 2,222,103 experimental wells, encompassing 17,063 CRISPR knockouts and 1,674 commercially available bioactive compounds tested at 8 concentrations each. Most CRISPR knockouts are represented by six unique guide RNAs targeting the gene.
RxRx3-core, outlined in figure 1, focuses on a smaller, more practical subset, including images from the 736 genes knockouts not blinded in RxRx3, all 1,674 compounds at all 8 concentrations, and relevant control wells. This curated subset retains the critical data necessary for benchmarking representation learning models while substantially reducing dataset size. The key statistics for RxRx3-core are summarized in Table 1.
Table 1: RxRx3-core summary statistics
Metric Count
Images (individual fluorescent channel JP2 files) 1,335,606
Wells 222,601
Plates sampled 1,744
Unique compounds 1,674
Unique compounds-concentration pairs 13,358
Unique CRISPR gene targets 736
Unique CRISPR guide RNAs 4,433
Unique CRISPR guide RNAs per gene target 6
# of well reps per CRISPR guide RNA 18
# of well reps per compounds-concentration pairs 4
2.1 Experimental Details
CRISPR Knockout Experiments: RxRx3-core includes data from gene001 to gene176 experiments in RxRx3. Each experiment includes 9 repeats of the same subset of perturbations but with randomized well layouts per plate. Each plate contains a combination of control and target guide wells, with only wells that pass quality control filters retained (Celik et al., 2024). Wells with annotations for 736 genes and a subset of control wells are included.
Compound Experiments: The dataset includes 1,674 compounds tested at 8 concentrations. Each compound experiment is typically repeated four times, with control compounds tested more frequently. As with CRISPR experiments, only wells that pass quality control filters (Celik et al., 2024) have been included in RxRx3-core.
Compressing RxRx3-core: The original RxRx3 dataset exceeded 100 TB and the majority of its well treatment metadata was blinded, making it challenging for widespread use within the research community. To address this limitation, we compressed the dataset to just 17.5 GB while preserving its utility for benchmarking. The data preparation steps applied are outlined in Table 2. These compression steps reduced the image set size for RxRx3-core from over 10 TB for the set of unblinded wells included in RxRx3-core to 17.5 GB without compromising the ability to distinguish representation learning methods effectively.
Table 2: Data processing steps to create RxRx3-core.
Data Preparation Step Compression Factor
Subset to unblinded wells in RxRx3 1/10
Included 512x512 center crops from original 2048x2048 images 1/16
Converted image format from uint16 to uint8 1/2
Applied JPEG 2000 compression using quality layers [80, 40, 20] 1/16
Accessibility of RxRx3-core: RxRx3-core is designed to be highly accessible to the machine learning community. The dataset is available on HuggingFace (huggingface.co/datasets/recursionpharma/rxrx3-core) and can be loaded directly using the HuggingFace datasets.load_dataset function. Additionally, embeddings generated by the OpenPhenom-S/16 (Kraus et al., 2024) model are provided on the dataset page. The dataset is also available on the Polaris platform (Wognum et al., 2024; 2025) for drug discovery (polarishub.io/datasets/recursion/rxrx3-core).
To facilitate benchmarking, we have released the benchmarking code in the EFAAR repository(github.com/recursionpharma/EFAAR_benchmarking) and on the Polaris platform (polarishub.io/benchmarks/recursion/rxrx-compound-gene-activity-benchmark), enabling a streamlined model evaluation and development.
3 Phenomic representation learning
On the RxRx3-core images, we evaluated a manual feature extraction method (CellProfiler) compared to three self-supervised vision transformer-based foundation models trained as masked autoencoders (MAE), as summarized in Table 3. Together, the four representation learning methods we evaluate provide a strong comparison between open and proprietary approaches for feature extraction in high-content phenotypic screening.
CellProfiler is a widely used open-source tool for cell segmentation and feature extraction. CellProfiler features for RxRx3-core were computed with version 2.2.0 (Kamentsky et al., 2011). Single cells were segmented after applying illumination correction and shape, intensity, and texture features were extracted from each cell, resulting in 952 dimensional profiles.
OpenPhenom-S/16 is a fully open-access model trained on publicly available datasets and is available on HuggingFace (huggingface.co/recursionpharma/OpenPhenom). It leverages a channel-agnostic MAE architecture (Kraus et al., 2024) with a 25 million parameter ViT-S/16 encoder backbone and 6 decoders (one for each channel). It was trained for 100 epochs on RxRx3 (Fay et al., 2023) and cpg0016 (CRISPR and ORF subsets from sources 4 and 13 of JUMP) (Chandrasekaran et al., 2023). Being channel-agnostic, the model tokenizes each image channel’s 16x16 pixel patches separately, yielding 1536 tokens for a 6-channel image at inference time and 384 unmasked tokens visible during its training with a 75% mask ratio.
Phenom-1 is a standard MAE (He et al., 2022) with a ViT-L/8 encoder backbone trained with a 75% mask ratio (Kraus et al., 2024). It is a proprietary model trained on over 3.5 billion image crops sampled from the RPI-93M dataset, which contains 93 million unique wells across a large variety of perturbations and cell types. Its 8x8x6 pixel patch size means that each image crop is tokenized to 1024 tokens, and during training the encoder processes 256 tokens to predict the rest.
Phenom-2 is a standard MAE with a ViT-G/8 encoder backbone trained with a 75% mask ratio (Kenyon-Dean et al., 2024). We take its embeddings from the trimmed layer, where linear probes found that the best representations from this model were at the intermediate 38th layer of the encoder, rather than the final 48th layer. This proprietary model was trained on over 8 billion microscopy image crops sampled from Phenoprints-16M, a specially curated dataset of 16 million wells with a diverse set of meaningful perturbations. Both Phenom-1 and Phenom-2 include non-public 6-channel brightfield and CellPainting images in their training datasets.
Table 3: Representation learning methods examined in this work.
Model Training Data Architecture Parameters
CellProfiler - - -
OpenPhenom-S/16 RxRx3 + cpg0016 CA-MAE ViT-S/16 25M
Phenom-1 RPI-93M MAE ViT-L/8 307M
Phenom-2 Phenoprints-16M MAE ViT-G/8 1,860M
4 Results
RxRx3-core enables the computation of benchmarks that evaluate the quality of perturbation embeddings in capturing biological signals across key tasks: distinguishing perturbations from control populations and zero-shot retrieval of drug-target interactions. Before calculating these benchmarks, crop aggregation and batch alignment techniques are applied to phenomics embeddings from representation learning models. First, the four tiled embeddings from every 256x256x6 crop of each 512x512x6 microscopy image are mean-aggregated. Next, a crucial batch alignment step is used to center the latent space on control samples and align experimental batches, ensuring comparability across conditions (Celik et al., 2024).
We use a PCA-based alignment technique, PCA-CS (Principal Component Analysis with Centering and Scaling), to achieve this. First, a PCA transformation is fitted on all control samples (without reducing dimensionality) and applied to all data points. Next, for each experimental batch, embeddings are centered and scaled relative to the corresponding batch controls. These pre-processing steps standardize the embeddings and minimize batch effects. Following alignment, we compute two benchmarks to assess the embeddings’ information content.
4.1 Perturbation signal benchmark
The energy distance (Rizzo & Székely, 2016) quantifies the separation between the distribution of replicate embeddings for a perturbation and that of negative controls, effectively serving as a measure of perturbation effect size in a high-dimensional space. To evaluate this, we calculate the distance between the distribution of embeddings for the replicate perturbation embeddings and the embeddings of the control units, employing statistical methods based on energy metrics.
Assuming access to two sets of embeddings $\mathbf{x}_{1},\dots,\mathbf{x}_{n_{1}}$ (representing query perturbation units) and $\mathbf{y}_{1},\dots,\mathbf{y}_{n_{2}}$ (representing negative control units), the energy distance is defined as
$\text{energy}_{g}=\frac{2}{n_{1}n_{2}}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\|\bm{x}_{i}-\bm{y}_{j}\|-\frac{1}{n_{1}^{2}}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{1}}\|\bm{x}_{i}-\bm{x}_{j}\|-\frac{1}{n_{2}^{2}}\sum_{i=1}^{n_{2}}\sum_{j=1}^{n_{2}}\|\bm{y}_{i}-\bm{y}_{j}\|.$
This metric equals zero when the two distributions are identical and increases with greater divergence between them. The benchmark assumes that better representation learning methods will embed more query perturbations farther from control replicates. In this study, we report the mean energy distance (± MAD) across all perturbations, alongside corresponding z-scores, to indicate improvement over a random baseline. In Table 4 we present the results for the perturbation magnitude benchmark using CellProfiler and ViT baselines described above.
4.2 Drug-target interaction benchmark
In addition to reporting perturbation signal magnitude, we also present a benchmark to assess how well representation learning models can relate CRISPR knockouts and compound perturbations in HCS datasets. We evaluate the benchmark for zero-shot prediction of drug-target interactions by analyzing cosine similarities between their embeddings (Figure 2). The compound-gene interaction annotations used in the benchmark were curated from PubChem, Guide to Pharmacology, WIPO, D3R, BindingDB, US Patents, and ChEMBL (Liu et al., 2007; Zdrazil et al., 2024; Harding et al., 2024). To ensure high confidence in our benchmark ground truth, we include only EC50 (IC50) values and avoid mixing measurement types from different assays. These values are aggregated by the median across experiments with the same compound-target pair. A subset of the resulting drug-target interactions, focusing on key genes, is summarized in Table 7. This approach assesses whether a model ranks the known target genes of each compound higher than a randomly selected set of genes from the ground truth dataset based on cosine similarity. To quantify confidence, we use the absolute value of the cosine similarity as a proxy, analogous to the probability score in a classifier. For each compound, we calculate the Area Under the ROC Curve (AUC) and average precision as performance metrics.
Figure 2: Mean average precision performance on RxRx3-core public benchmark in predicting drug-target interactions, across all compound concentrations with error bars for 100 runs of the benchmark with different random seeds (Table 4).
The final results are summarized by reporting the median AUC and average precision across all compounds, providing a comparison against a random baseline. Given that this precision benchmark requires randomly sampling negative interactions, we report results for 100 random negative sampling seeds of the benchmark with mean and standard deviation. This allows robust comparison against performance of a random baseline via z-score:
$z=\frac{\bar{x}_{\text{model}}-\bar{x}_{\text{baseline}}}{\sqrt{s_{\text{model}}^{2}+s_{\text{baseline}}^{2}}}$
Table 4 highlights the detailed results along the max axis, which captures the strongest potential interaction for each compound irrespective of its experimental concentration. We note that, while CellProfiler does create salient representations of samples which are more distinguished from controls (Energy), this performance does not necessarily correlate to the primary compound-gene benchmark. These results provide evidence that deep learning features from the Phenom models are better at capturing compound-gene activity in this image data than the CellProfiler features. Overall, Phenom-2’s strong performance on each task demonstrates the importance of scaling self-supervised learning to yield biologically meaningful representations, and furthermore motivates continued development of deep learning techniques to continue improving drug discovery.
Table 4: Performance on the public RxRx3-core drug-target interaction and RxRx3-core perturbation magnitude benchmark, measuring mean average precision ($\pm$ STD over 100 random seeds benchmarking with different negative samples) in predicting compound activity against target genes, and mean energy distance ($\pm$ MAD over all perturbations) separating perturbation embeddings from controls with corresponding z-scores of improvement over a random baseline.
Method Average Precision Comp. z-score ↑ Energy dist. Energy z-score ↑ Random embeddings $0.214\pm 0.003$ 0.00 $0.728\pm 0.05$ 0.00 Cell Profiler $0.276\pm 0.018$ 3.34 $6.245\pm 1.32$ 4.165 OpenPhenom-S/16 $0.277\pm 0.016$ 3.89 $3.315\pm 0.78$ 3.310 Phenom-1 $0.290\pm 0.017$ 4.35 $4.848\pm 1.07$ 3.836 Phenom-2 $\textbf{0.307}\pm 0.015$ 6.04 $\mathbf{6.636}\pm 1.33$ 4.435
5 Conclusion
We introduce RxRx3-core, a compact and accessible dataset with an associated set of benchmarks designed to evaluate representation learning models for high-content screening microscopy data against zero-shot drug-target interaction prediction. At just 18GB, RxRx3-core is available on HuggingFace and Polaris, making it practical for widespread use. Our analysis includes baseline and foundation model performance across perturbation signal magnitude and zero-shot drug-target interaction prediction benchmarks. The results highlight the advantages of self-supervised vision transformer models, particularly as they scale to multi-billion parameter regimes, in capturing biologically meaningful representations. We hope that RxRx3-core and its benchmarks serve as a valuable resource for the machine learning community, fostering innovation in representation learning approaches and advancing the analysis of large-scale microscopy datasets.
Meaningfulness Statement
A meaningful representation of life characterizes relevant aspects of the vital functions of a sample. This can be measured by benchmarking representations of experimental data to real-world known relationships. We evaluate different representation learning methods on images of HUVEC cells in our newly released public dataset, RxRx3-core, to determine how those cells are phenotypically impacted by genetic and chemical perturbations. The most advanced self-supervised model, Phenom-2 (Kenyon-Dean et al., 2024), creates the most meaningful representations in this context. We believe that this work will enable development of new methods for connecting fundamental genetic functions with compound activity for the purposes of drug discovery.
Acknowledgments
We thank M. McKeithen, S. Moore, E. Hurst, and J. Fryer for their help computing and accessing CellProfiler features for this manuscript and the incredible Recursion lab and engineering teams for design and execution of experiments and storage and processing of data.
Appendix A Appendix
A.1 Gene-gene interaction benchmark
RxRx3-core can potentially be used to benchmark representation learning models against zero-shot gene-gene interaction prediction. However, we caution users from making conclusions based on evaluating this task on the RxRx3-core dataset given the relatively small number of CRISPR knockouts in the dataset (736). The task, previously described in Celik et al. (2024) and evaluated in Kraus et al. (2024) and Kenyon-Dean et al. (2024), leverages embeddings from representation learning models to build biological maps that can be used to uncover novel and meaningful interactions between perturbations. To assess the ability of a map to represent known biological relationships, we employ four annotation sources containing gene-gene interactions as benchmarks (Table 5).
Table 5: Description of datasets used for curating gene-gene interactions.
Dataset Description
CORUM Giurgiu et al. (2019) Gene clusters corresponding to protein complexes
HuMAP (Drew et al., 2017) Gene clusters representing protein complexes
Reactome (Gillespie et al., 2021) Protein-protein interactions based on curated pathways
StringDB (Szklarczyk et al., 2020) Proteins associations based on functional interactions
The core hypothesis is that a map capable of accurately capturing known relationships provides a strong representation of biological systems and is likely to reveal novel interactions. To evaluate this, closely following Celik et al. (2024) Multivariate known biological relationship recall benchmark, we compute recall for the most extreme 10% of pairwise relationships based on cosine similarity between embeddings. This includes both the top 5% (high similarity) and bottom 5% (low similarity) of the pairwise similarity distribution, as the latter may indicate opposing functions between genes. A random or uninformative map serves as a baseline, expected to achieve a recall of 10%.
To calculate recall, we first compute pairwise cosine similarities between the aggregated embeddings of all perturbed genes in RxRx3-core. Self-links, where a gene is compared to itself, are excluded because their cosine similarity is always one, which would bias the results. “Predicted links” are defined as relationships falling within the top or bottom 5% of the similarity distribution. Recall is then determined for each annotation source as the proportion of true predicted links relative to all possible links. Since datasets differ in the number of genes they include, we adjust recall calculations to consider only gene-gene interactions where both genes are present in the dataset. Therefore, recall results (Table 6) from the subset of 736 gene knockouts (KOs) in RxRx3-core are not directly comparable with other work (Kraus et al., 2024; Kenyon-Dean et al., 2024) which measures recall across a broader set of 17,000 genes.
To demonstrate how models rank on this metric when a larger set of gene-KOs are leveraged, we report the same gene-gene relationship recall benchmark (Celik et al., 2024) on the full RxRx3 dataset, including 17,063 gene-KOs, and on the JUMP dataset (cpg0016) (Chandrasekaran et al., 2023), including a subset of 7,976 gene-KOs, in Table 6. As shown in the table, the number of interactions evaluated grows significantly as more CRISPR knockouts are included in the datasets. For post-processing RxRx3-core and JUMP-CP embeddings, we use PCA with center-scaling for standard batch correction alignment. We observe that the MAEs perform better than the CellProfiler feature extraction baseline (Carpenter et al., 2006; Kamentsky et al., 2011) when a larger set of genes and gene-gene interactions are evaluated. For post-processing the full RxRx3 dataset, we correct for batch effects using Typical Variation Normalization (TVN) (Ando et al., 2017), and also correct for possible chromosome arm biases known to exist in CRISPR-Cas9 HCS data (Lazar et al., 2024). Notably, recall on JUMP-CP is considerably lower than on RxRx3 likely due to different assay protocols and more variance in the data.
Model CORUM HuMAP Reactome StringDB
RxRx3-core
736 genes  # relationships 1,209 958 569 1,737
Random embeddings .107 .111 .107 .115
CellProfiler .748 .813 .473 .652
OpenPhenom-S/16 .654 .709 .436 .579
Phenom-1 .721 .784 .402 .623
Phenom-2 .731 .810 .407 .647
RxRx3
17,063 genes  # relationships 29,459 47,699 14,493 43,707
CellProfiler .503 .354 .190 .414
OpenPhenom-S/16 .549 .374 .229 .429
Phenom-1 .609 .434 .251 .472
Phenom-2 .615 .437 .263 .492
JUMP-CP (cpg0016)
7,796 genes  # relationships 8,401 9,280 8,236 17,548
CellProfiler .219 .184 .131 .191
OpenPhenom-S/16 .233 .199 .154 .214
Phenom-1 .248 .208 .160 .226
Phenom-2 .264 .215 .165 .235
Table 6: Multivariate known biological relationship recall estimated @ 0.05-0.95 percentiles for benchmark databases (CORUM, hu.MAP, Reactome-PPI, and StringDB) over 3 random seeds sampling the null distribution (SD ±$\leq\pm 10^{-3}$). For RxRx3-core dataset the embeddings are batch-corrected using a PCA-CS aligner. For RxRx3 dataset, the embeddings are transformed with a TVN aligner Ando et al. (2017) followed by a chromosome arm bias correction Lazar et al. (2024). For JUMP-CP (cpg0016) dataset we used a PCA-CS aligner.
Gene Compounds
ACE Moexipril, Losartan, Quinapril, Enalapril, Benazepril, Sitagliptin phosphate monohydrate, bucillamine, zofenopril-calcium
AR Flutamide, Enzalutamide, Liothyronine, Indomethacin, Mifepristone, ODM-201, ARN-509, Eugenol, Nilutamide, Eplerenone, Pamoic acid disodium salt, Dihydrotestosterone(DHT), salicylamide, bicalutamide
BRAF Gefitinib, Imatinib, Sorafenib, Erlotinib, Vemurafenib, Dabrafenib, BIRB 796, PD0325901, Regorafenib, RepSox, Niclosamide, CHEMBL410295, MLN 2480
EGFR Adrucil, Flavopiridol, Gefitinib, Lapatinib, BMS-777607, Phloretin, Crizotinib, BMS-536924, Quercetin, Tivozanib, Imatinib, Foretinib, Sorafenib, Erlotinib, Vemurafenib, OSI-906, Indomethacin, PD168393, Icotinib, BIRB 796, LDK378, NVP-AEW541, Pexidartinib, Ketoprofen, Vorinostat, Aspirin, AZD9291, Ibrutinib, Cabozantinib, Vandetanib, Ponatinib, LY2228820, Entrectinib, Dasatinib, Bosutinib, Niclosamide, Olmutinib, Acalabrutinib, Curcumin, brigatinib
ESR1 Phloretin, Nabumetone, Raloxifene, Bazedoxifene, Letrozole, Mifepristone, Hexachlorophene, Molsidomine, Nordihydroguaiaretic acid, Resorcinol, Eplerenone
HDAC1 Atorvastatin, Divalproex, Fluconazole, Bendamustine, Lapatinib, Imatinib, Celecoxib, Daunorubicin, Givinostat, Phenylbutyrate, PCI-24781, Vorinostat, AZD9291, Bortezomib, Vandetanib, LMK-235, BIIB021, Colchicine, Bestatin, Curcumin, lovastatin
MTOR BGT226, GSK2126458, AZD2014, CAL-101, GDC-0349, BYL719, Dasatinib, PF-04691502, INK 128, Gedatolisib, Torin 1, LY3023414, GDC-0084, Dactolisib (BEZ235)
PTGS2 Acetaminophen, Diclofenac, Naproxen, Quercetin, Vortioxetine, Celecoxib, Erlotinib, Meloxicam, Mesalamine, Indomethacin, Nimesulide, Lumiracoxib, Zileuton, Ketoprofen, Valdecoxib, Aspirin, Piroxicam, Salicylate, Eugenol, Nordihydroguaiaretic acid, Parecoxib, Loxoprofen, Curcumin, Etoricoxib, Flurizan, ibuprofen-(s), meclofenamic-acid, flurbiprofen-(S)-(+), carprofen, ketorolac, deracoxib, tepoxalin, firocoxib
SRC AMG-458, Flavopiridol, Crizotinib, Quercetin, Imatinib, Nilotinib, PD168393, BIRB 796, NVP-AEW541, Enzastaurin, Ibrutinib, Cabozantinib, Vandetanib, Ponatinib, CH5183284, PRT062607, Dasatinib, Bosutinib, Mebendazole, Niclosamide, NVP-TAE 684, CHEMBL410295, AZM475271, Acalabrutinib, fosfosal
TP53 BYL719
Table 7: Known gene-compound associations used for benchmarking, showing FDA-approved drugs with their validated molecular targets. The exemplary subset includes major therapeutic targets such as kinases (EGFR, BRAF), nuclear receptors (AR, ESR1), and enzymes (ACE, PTGS2).